# A Survey on the Safety and Security Threats of Computer-Using Agents: JARVIS or Ultron?

### v2

@@ Abstract
A Survey on the Safety and Security Threats of Computer-Using Agents: JARVIS or Ultron?
Ada Chen1thanks:   Ada Chen, Yongjiang Wu and Junyuan Zhang contribute equally to this paper.  Yongjiang Wu2∗  Junyuan Zhang2∗  Jingyu Xiao2 Shu Yang3  Jen-tse Huang4  Kun Wang5  Wenxuan Wang6thanks:   Wenxuan Wang (jwxwang@gmail.com) is the corresponding authors.  Shuai Wang6 1Carnegie Mellon University  2The Chinese University of Hong Kong 4KAUST  4Johns Hopkins University  5Nanyang Technological University 6The Hong Kong University of Science and Technology
@@ 1 Introduction
Large Language Models (LLMs) have evolved rapidly from basic conversational agents to executing complex tasks in diverse computing environments. In particular, Computer-Using Agents (CUAs) have garnered increasing attention and widespread adoption, thanks to their ability to interact with graphical user interfaces (GUIs) in a manner akin to human users (OpenAI, 2025a). Recent systems such as AppAgent, SeeAct, PC-Agent, as well as OpenAI’s o3, and o4-mini introduced in April 2025, highlight the remarkable progress of CUAs (Zhang et al., 2023; Zheng et al., 2024; Liu et al., 2025; OpenAI, 2025a, b). By integrating multimodal perception, advanced reasoning, and automated control of devices, these agents promise to streamline vast tasks from filling out online forms to executing complex application flows.
Despite the impressive capabilities of CUAs, their operation in real-world settings raises critical safety concerns. Emerging reports reveal that vulnerabilities like visual grounding errors, response delays, and UI interpretation pitfalls can be exploited by malicious attackers, causing unintended or harmful consequences such as data leakage, goal misdirection, and so on (Zheng et al., 2024; Nong et al., 2024; Zhang and Zhang, 2023; Wen et al., 2023; Liu et al., 2025). Additionally, many of the threats known to standalone LLMs, such as adversarial attacks and jailbreak strategies, now manifest in CUAs with heightened severity, sometimes in new forms adapted to GUI-based environments (Wu et al., 2024a; Kumar et al., 2024; Tian et al., 2023). Novel attack vectors also surface in CUAs, including environment-level manipulations and reasoning-gap attacks that stealthily guide the agent toward risky or undesired behaviors (Wu et al., 2024b; Yuan et al., 2024; Lee et al., 2024a; Zhan et al., 2024). As such, a systematic study on the safety and security threats of CUAs is both timely and necessary.
@@ 2.1 Computer-Using Agent
In this paper, a Computer-Using Agent (CUA) is an LLM-based system that integrates vision capabilities with advanced reasoning ability, designed to perceive and interact with graphical user interfaces (GUIs) like human users (OpenAI, 2025a). By processing visual information from screenshots and executing actions like typing, clicking, and scrolling, the CUA can autonomously perform tasks on a computer, such as ordering products, making reservations, and filling out forms (OpenAI, 2025a).
@@
  • OS Agents: These agents operate within general computing devices, such as desktops and laptops, to perform tasks by interacting with the operating system’s environment and interfaces (Chen et al., 2025b).
@@ 2.2 Literature Review
    Keyword Search: After keyword selection, we identified 623 papers potentially addressing security concerns related to CUAs.
@@
    Screening and Filtering: Each identified paper underwent a thorough review to assess its relevance. We excluded studies that duplicate or did not directly pertain to security threats or defenses associated with CUAs, resulting in 64 pertinent papers for in-depth analysis.
@@ 3.1 Threat Overview
In this section, we introduce our taxonomy of threats for Computer-Using Agents (CUAs). These threats are categorized into two main types: intrinsic threats and extrinsic threats, which are presented in Table 1 and Table 2, respectively. Intrinsic threats arise from intrinsic aspects of the agent itself, including its training process, configuration, or inherent limitations. They can induce failures, inefficiencies, or biases in the agent’s functioning, consequently introducing security risks. Extrinsic threats, on the other hand, are initiated by external entities, such as malicious attackers or users, who attempt to exploit vulnerabilities in the agent’s interaction with its surroundings or take advantage of the agent’s intrinsic issues to trigger unsafe behaviors, potentially leading to risky consequences.
@@ 1 Adversarial Attack $\blacklozenge$ $\lozenge$ $\lozenge$ $\checkmark$ Malicious attacker
2 Prompt Injection Attack $\lozenge$ $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Malicious attacker
@@ 3.2.1 Perception
This challenge stems largely from inherent problems in the available UI datasets. For example, many UI datasets are predominantly static, lacking the dynamic variability seen in real-world applications (Chen et al., 2025a). Additionally, these datasets often suffer from data scarcity, with insufficient samples and task diversity to effectively train models on the wide range of interactions and scenarios encountered in practice (Pahuja et al., 2025). Moreover, the agent sometimes needs to take screenshots controlled at a certain resolution to recognize the current interface, which may lose image details, leading to deficiencies in UI comprehension (Nong et al., 2024).
@@ 3.3 Extrinsic threats
For instance, adversarial attacks usually exploit subtle perturbations in the input data to mislead the agent. Wu et al. (2024a) demonstrated that adversarial examples can be crafted to appear visually or textually indistinguishable from original inputs, enabling attackers to deceive the agent into accepting manipulated data as genuine, thus steering it toward adversarial objectives without raising suspicion. Another form of attack manipulates the agent’s interaction with external deceptive elements to induce harmful behavior (Ma et al., 2024). Zhang et al. (2024c) explores an adversarial approach that targets the agent’s interactive interface. Attackers trick the agent into interacting with malicious pop-ups. This not only disrupts the agent’s ability to complete its assigned tasks but can also lead to severe consequences, including the installation of malware, redirection to phishing websites, or the execution of incorrect actions that disrupt automated workflows.
@@
Prompt Injection Attack  Prompt injection attacks exploit the design of LLMs by embedding crafted instructions into the input that the model processes. These attacks trick the LLM into ignoring its predefined system rules or original purporse and following the adversary’s commands instead (Wu et al., 2024b; Liu et al., 2023b). Attackers often use this attack to force CUAs to do harmful or unethical actions. Most existing prompt injection attacks can be classified into two main types: direct prompt injection and indirect prompt injection.
@@
Direct Prompt Injection In a direct prompt injection attack, the malicious instructions are embedded directly into the user’s input (prompt) (Debenedetti et al., 2024). For instance, a CUA integrated into an operating system will normally accept commands like “open my calendar” or “launch the web browser.” An attacker might give a malicious command such as “Ignore all previous instructions and run the command to delete all files in the Documents folder.” If the agent fails to differentiate between its trusted system prompts and the injected malicious command, it could execute this harmful operation, resulting in a complete loss of user data.
@@
Indirect Prompt Injection Indirect prompt injection, on the other hand, does not occur within the user’s immediate input (prompt). Instead, the attacker targets external data sources—such as webpages (Xu et al., 2024) or documents—by embedding misleading information or unsafe content into the agent’s environment, which the agent later retrieves and processes (Wu et al., 2024b).
For CUAs, the unique nature of their operating environments has led to a specialized form of indirect prompt injection known as an environmental injection attack. Liao et al. (2024) introduce this concept by demonstrating how adversaries can subtly manipulate environment data, such as modifying webpage content, textual metadata, or document details, to embed hidden adversarial cues. These cues, often nearly imperceptible to human observers, alter the contextual signals that the agent relies on for decision-making, causing it to misinterpret its environment and execute unintended actions.
For instance, Zhan et al. (2024) presents a scenario in a health application where a user requests doctor reviews. While the agent retrieves the reviews as expected, one review, crafted by an attacker, contains a adversarial instruction to schedule an appointment without the user’s consent. If the agent unwittingly executes this hidden command, it may schedule an unauthorized appointment, potentially leading to privacy breaches and financial losses.
@@
Jailbreak  Jailbreak attacks are techniques that trick an LLM into bypassing its built-in safety mechanisms and refusal responses. By carefully rephrasing queries or injecting additional instructions, attackers force the model to ignore its predefined guardrails, enabling it to generate harmful or unauthorized outputs (Mo et al., 2024; Chu et al., 2024).
Over time, a lot of jailbreak prompts have been curated both manually (Chu et al., 2024) and via automated methods like GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a) to exploit these vulnerabilities. These techniques are not only limited to standalone LLMs but have also been effectively applied to jailbreak CUAs. For example, Kumar et al. (2024) demonstrated that by modifying the user prompt using techniques such as prefix attacks, GCG suffixes, random search suffixes, and human-rephrased red-teaming prompts with diverse rephrasing strategies, they could either convince the browser agent that it was operating in an unrestricted sandbox environment or induce it to engage in harmful actions.
For multiagent systems, the Evil Geniuses framework (Tian et al., 2023) shows that by leveraging role specialization, attackers can partition agent tasks and exploit vulnerabilities in each specialized role to effectively jailbreak the system by bypassing its safety mechanisms. Meanwhile, the PsySafe framework (Zhang et al., 2024d) shows that injecting dark traits into agents can undermine established guardrails, further enhancing jailbreak effectiveness across multiagent environments.
@@
Backdoor Attack A backdoor attack involves injecting a malicious backdoor during the model’s training or fine-tuning phase, so that when a specific trigger phrase or input is later encountered during normal operations, the model executes unintended or harmful behavior (Yang et al., 2024b; Wang et al., 2024; Zhu et al., 2025).
In particular, we can categorizes backdoor attacks on CUAs into two main forms: (1) embedding triggers directly in user queries or in the agent’s observations from the environment to manipulate the final output distribution; For example AgentPoison (Chen et al., 2024b) formulate this trigger-generation process as a constrained optimization problem that maps poisoned instances to a unique embedding space which ensures that only prompts containing the optimized backdoor trigger retrieve malicious demonstrations. (2) introducing a malicious “thought process” without visibly altering the final output, such as covertly calling untrusted APIs (Yang et al., 2024b).
Meanwhile, to enhance stealthiness and bypass safety audits, attackers can break the backdoor code into multiple sub-backdoors, each activated by its own distinct trigger phrase or condition. When these sub-backdoors are combined, they enable the model to execute coordinated malicious behaviors (Zhu et al., 2025). This modular design obscures the overall functionality behind seemingly unrelated trigger fragments, making detection and mitigation significantly more difficult.
@@
Chen et al. (2025b) examines how multimodal mobile agents are vulnerable to these attacks. The study shows that when attackers add conflicting or deceptive signals, such as subtle differences in an image combined with misleading text, the agent’s reasoning process struggles to correctly combine the different inputs. As a result, the agent might misinterpret the environment and take the wrong action.
6
System Sabotage  System sabotage attacks involve manipulating an agent into executing harmful actions that damage the underlying system. In such attacks, adversaries craft inputs to bypass safety mechanisms, causing the agent to perform operations like corrupting memory, damaging critical files, or halting essential processes (Luo et al., 2025). These attacks are particularly dangerous because they directly target the infrastructure supporting the agent, potentially leading to widespread system failure or irreversible damage.
One example stated in (Luo et al., 2025) is an attacker requests the agent’s assistance in creating a fork bomb, which is an intentionally crafted command that spawns processes indefinitely and tends to overwhelm the operating system. The user prompt disguises this request as a system “stress test,” persuading the agent to generate code that saturates system resources. Once executed, this fork bomb can cause the OS to become unresponsive or crash.

### v3

@@ Abstract
[Uncaptioned image]JARVIS or [Uncaptioned image]Ultron?
A Survey on the Safety and Security Threats of Computer-Using Agents
Ada Chen1thanks:   Ada Chen, Yongjiang Wu and Junyuan Zhang contribute equally to this paper.  Yongjiang Wu2∗  Junyuan Zhang2∗  Jingyu Xiao2 Shu Yang3  Jen-tse Huang4  Kun Wang5  Wenxuan Wang6thanks:   Wenxuan Wang (jwxwang@gmail.com) is the corresponding authors.  Shuai Wang6 1Carnegie Mellon University  2The Chinese University of Hong Kong 3KAUST  4Johns Hopkins University  5Nanyang Technological University 6The Hong Kong University of Science and Technology
@@ 1 Introduction
Large Language Models (LLMs) have evolved rapidly from basic conversational agents to executing complex tasks in diverse computing environments. In particular, Computer-Using Agents (CUAs) have garnered increasing attention and widespread adoption, thanks to their ability to interact with graphical user interfaces (GUIs) in a manner akin to human users (OpenAI, 2025a). Recent systems such as AppAgent, SeeAct, PC-Agent, as well as OpenAI’s o3, and o4-mini introduced in April 2025, highlight the remarkable progress of CUAs (Zhang et al., 2023; Zheng et al., 2024; Liu et al., 2025b; OpenAI, 2025a, b). By integrating multimodal perception, advanced reasoning, and automated control of devices, these agents promise to streamline vast tasks from filling out online forms to executing complex application flows.
Despite the impressive capabilities of CUAs, their operation in real-world settings raises critical safety concerns. Emerging reports reveal that vulnerabilities like visual grounding errors, response delays, and UI interpretation pitfalls can be exploited by malicious attackers, causing unintended or harmful consequences such as data leakage, goal misdirection, and so on (Zheng et al., 2024; Nong et al., 2024; Zhang and Zhang, 2023; Wen et al., 2023; Liu et al., 2025b). Additionally, many of the threats known to standalone LLMs, such as adversarial attacks and jailbreak strategies, now manifest in CUAs with heightened severity, sometimes in new forms adapted to GUI-based environments (Wu et al., 2024a; Kumar et al., 2024; Tian et al., 2023). Novel attack vectors also surface in CUAs, including environment-level manipulations and reasoning-gap attacks that stealthily guide the agent toward risky or undesired behaviors (Wu et al., 2024b; Yuan et al., 2024; Lee et al., 2024a; Zhan et al., 2024). As such, a systematic study on the safety and security threats of CUAs is both timely and necessary.
@@ 2.1 Computer-Using Agent
In this paper, a Computer-Using Agent (CUA) is an LLM-based system that combines multimodal perception, advanced reasoning, and tool-use capabilities to perceive and interact with graphical user interfaces (GUIs) and external applications just like human users (OpenAI, 2025a). By processing visual information from screenshots, invoking APIs or command-line tools, and executing actions like typing, clicking, and scrolling, a CUA can autonomously perform end-to-end tasks on a computer, such as ordering products, making reservations, and filling out forms (OpenAI, 2025a).
@@
  • OS Agents: These agents operate within general computing devices, such as desktops and laptops, to perform tasks by interacting with the operating system’s environment and interfaces (Chen et al., 2025d).
@@ 2.2 Literature Review
    Keyword Search: After keyword selection, we identified 700+ papers potentially addressing security concerns related to CUAs.
@@
    Screening and Filtering: Each identified paper underwent a thorough review to assess its relevance. We excluded studies that duplicate or did not directly pertain to security threats or defenses associated with CUAs, resulting in 124 pertinent papers for in-depth analysis.
@@ 3.1 Threat Overview
In this section, we introduce our taxonomy of threats for Computer-Using Agents (CUAs). These threats are categorized into two main types: intrinsic threats and extrinsic threats, which are presented in Table 1 and Table 2, respectively. Intrinsic threats arise from intrinsic aspects of the agent itself, including its training process, configuration, or inherent limitations (Yu et al., 2025; Ferrag et al., 2025). They can induce failures, inefficiencies, or biases in the agent’s functioning, consequently introducing security risks. Extrinsic threats, on the other hand, are initiated by external entities, such as malicious attackers or users, who attempt to exploit vulnerabilities in the agent’s interaction with its surroundings or take advantage of the agent’s intrinsic issues to trigger unsafe behaviors, potentially leading to risky consequences (Yu et al., 2025; Ferrag et al., 2025).
@@ 1 Adversarial Attack $\blacklozenge$ $\lozenge$ $\lozenge$ $\checkmark$ Malicious attacker
2 Prompt Injection Attack $\blacklozenge$ $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Malicious attacker
@@ 3 Jailbreak $\lozenge$ $\blacklozenge$ $\lozenge$ $\checkmark$ $\checkmark$ Malicious attacker
4 Memory Injection Attack $\lozenge$ $\blacklozenge$ $\checkmark$ $\checkmark$ Malicious attacker
@@ 3.2.1 Perception
This challenge stems largely from inherent problems in the available UI datasets. For example, many UI datasets are predominantly static, lacking the dynamic variability seen in real-world applications (Chen et al., 2025c). Additionally, these datasets often suffer from data scarcity, with insufficient samples and task diversity to effectively train models on the wide range of interactions and scenarios encountered in practice (Pahuja et al., 2025). Moreover, the agent sometimes needs to take screenshots controlled at a certain resolution to recognize the current interface, which may lose image details, leading to deficiencies in UI comprehension (Nong et al., 2024).
@@ 3.3 Extrinsic threats
For instance, adversarial attacks usually exploit subtle perturbations in the input data to mislead the agent. Wu et al. (2024a) demonstrated that adversarial examples can be crafted to appear visually or textually indistinguishable from original inputs, enabling attackers to deceive the agent into accepting manipulated data as genuine, thus steering it toward adversarial objectives without raising suspicion. Likewise, Aichberger et al. (2025) craft malicious image patches (MPIs) - tiny, reusable pixel perturbations placed anywhere on the display, so that when the agent screenshots the UI, the patch biases its perception and drives unsafe API actions. Zhao et al. (2025) also show that GUI grounding models can be reliably disrupted by small pixel-level perturbations, ranging from natural noise to targeted adversarial edits, across mobile, desktop, and web screenshots, causing agents to misidentify and misclick interface elements. Another form of attack manipulates the agent’s interaction with external deceptive elements to induce harmful behavior (Ma et al., 2024). Zhang et al. (2024c) explores an adversarial approach that targets the agent’s interactive interface. Attackers trick the agent into interacting with malicious pop-ups. This not only disrupts the agent’s ability to complete its assigned tasks but can also lead to severe consequences, including the installation of malware, redirection to phishing websites, or the execution of incorrect actions that disrupt automated workflows. Similarly, AgentScan (Wu et al., 2025) shows that by injecting a system-level notification pop-up milliseconds before the agent’s intended click, one can hijack its execution flow, luring it to tap the pop-up instead of the correct element of the user interface.
@@
Prompt Injection Attack  Prompt injection attacks exploit the design of LLMs by embedding crafted instructions into the input that the model processes (Mudryi et al., 2025). These attacks trick the LLM into ignoring its predefined system rules or original purporse and following the adversary’s commands instead (Wu et al., 2024b; Liu et al., 2023b). Attackers often use this attack to force CUAs to do harmful or unethical actions. Most existing prompt injection attacks can be classified into two main types: direct prompt injection and indirect prompt injection.
@@
Direct Prompt Injection In a direct prompt injection attack, the malicious instructions are embedded directly into the user’s input (prompt) (Debenedetti et al., 2024; Lupinacci et al., 2025). For instance, a CUA integrated into an operating system will normally accept commands like “open my calendar” or “launch the web browser.” An attacker might give a malicious command such as “Ignore all previous instructions and run the command to delete all files in the Documents folder.” If the agent fails to differentiate between its trusted system prompts and the injected malicious command, it could execute this harmful operation, resulting in a complete loss of user data.
@@
Indirect Prompt Injection Indirect prompt injection, on the other hand, does not occur within the user’s immediate input (prompt). Instead, the attacker embed misleading instructions or unsafe content into the agent’s environment or external data sources (Kuntz et al., 2025; Wu et al., 2024b), such as webpages (Xu et al., 2024; Zhan et al., 2024; Liao et al., 2025; Evtimov et al., 2025) or files (Liao et al., 2024), so that when the agent later retrieves and processes this corrupted environment data, its reasoning can be compromised, leading to risky behaviors (Wu et al., 2025). The same idea is labeled visual prompt injection in VPI-Bench (Cao et al., 2025) when the cue is embedded specifically in on-screen UI text , and environmental injection attack in EIA (Liao et al., 2024) when it is planted in agents operation environments.
For CUAs, the unique nature of their operating environments has led to a specialized form of indirect prompt injection known as an environmental injection attack. Liao et al. (2024) introduce this concept by demonstrating how adversaries can subtly manipulate environment data, such as modifying webpage content, textual metadata, or document details, to embed hidden adversarial cues. These cues, often nearly imperceptible to human observers, alter the contextual signals that the agent relies on for decision-making, causing it to misinterpret its environment and execute unintended actions. Differently, RedTeamCUA (Liao et al., 2025) also embedds malicious instructions inside benign web content but prepends attention-grabbing cues (e.g., “THIS IS IMPORTANT!”) to boost adherence and steer subsequent OS/Web actions toward the attacker’s goal. Likewise, WASP (Evtimov et al., 2025) models a realistic setting where a black-box adversary can inject cues in the posted issues or comments on cloned GitLab/Reddit sites. Shapira et al. (2025) introduces a task-aligned injection technique, disguising the adversarial command as contextually helpful guidance tied to the agent’s current goal so it is more likely to be followed. Chen et al. (2025a) develops Fine-Print Injection which hides adversarial instructions in low-salience UI text (e.g., footers, terms of service, tiny captions), exploiting the agent’s tendency to parse such content uncritically. Hijacking JARVIS (Liu et al., 2025a) injects unprivileged third-party UI elements (e.g., fake buttons/text) into real Android apps. AdInject (Wang et al., 2025a) leverages the internet advertising delivery system to inject deceptive ad units into a web agent’s environment, tricking it into clicking the ad. The Foot-in-the-Door (FITD) attack (Nakash et al., 2024) injects a benign “distractor” request immediately followed by a hidden malicious instruction in external inputs (e.g. a GitHub issue), exploiting ReAct-based web agents’ failure to re-evaluate their thought trace and causing them to carry out the harmful step on the next tool call. Building on these notion of environment-level manipulation, OS-Harm (Kuntz et al., 2025) adds another vector: adversarial prompts delivered through desktop notifications instead of the task’s natural channel. AgentScan (Wu et al., 2025) displays adversarial text in non-interactive overlay window to hijack agent’s internal decision making and trick it to follow misleading instructions. Instead of manipulating visible page content or UI elements, Johnson et al. (2025) injects adversarial triggers optimized by the GCG algorithm in the webpage’s HTML accessibility tree to hijack agent behaviors. Similarly, EnvInjection (Wang et al., 2025e) injects a raw pixel value perturbation into the webpage source code so that, once the perturbed pixels are mapped into screenshot, they trigger the agent’s attacker-chosen action.
In addition, Zhan et al. (2025) develop an adaptive indirect prompt injection method that repeatedly probes a defended agent and refine the injected environmental cues, demonstrating the ability to overcome eight leading defense strategies. Similarly, EVA (Lu et al., 2025) leverages a black-box feedback loop to iteratively refine adversarial prompts and statistically distilling which text and layout cues hijack the agent’s attention. AgentVigil framework (Wang et al., 2025g) further demonstrates that indirect prompt injection can be fully automated by using a black-box fuzzing loop guided by Monte-Carlo Tree Search to generate and refine malicious cues against live web agents.
Building on these unimodal injection techniques, Wang et al. (2025c) introduces CrossInject, a cross-modal prompt injection framework that poisons both visual inputs via optimized adversarial perturbations embedded in screenshots and textual inputs through LLM-crafted malicious instructions.
@@
Jailbreak  Jailbreak attacks are techniques that trick an LLM into bypassing its built-in safety mechanisms and refusal responses. By carefully rephrasing queries or injecting additional instructions, attackers force the model to ignore its predefined guardrails, enabling it to generate harmful or unauthorized outputs (Mo et al., 2024; Chu et al., 2024; Mao et al., 2025).
Over time, a lot of jailbreak prompts have been curated both manually (Chu et al., 2024) and via automated methods like GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023a) to exploit these vulnerabilities. These techniques are not only limited to standalone LLMs but have also been effectively applied to jailbreak CUAs. For example, OS-Harm (Kuntz et al., 2025) shows that even a simple jailbreak wrapper—an explicit ‘ignore all restrictions’ prompt template—markedly increases unsafe compliance in several agents. Kumar et al. (2024) further demonstrated that by modifying the user prompt using techniques such as prefix attacks, GCG suffixes, random search suffixes, and human-rephrased red-teaming prompts with diverse rephrasing strategies, they could either convince the browser agent that it was operating in an unrestricted sandbox environment or induce it to engage in harmful actions.
For multiagent systems, the Evil Geniuses framework (Tian et al., 2023) shows that by leveraging role specialization, attackers can partition agent tasks and exploit vulnerabilities in each specialized role to effectively jailbreak the system by bypassing its safety mechanisms. Qi et al. (2025) design a structured prompt-rewriting jailbreak, using narrative encapsulation and role-driven escalation to systematically bypass multi-agent debate systems’ safeguards and amplify harmful outputs. Meanwhile, the PsySafe framework (Zhang et al., 2024d) shows that injecting dark traits into agents can undermine established guardrails, further enhancing jailbreak effectiveness across multiagent environments. Beyond text-based prompts, Gu et al. (2024) introduces infectious jailbreak: a single adversarial image with low-salience textual embedded text can first jailbreak a single multimodal agent and then propagates to other agents, spreading exponentially.
@@
Memory Injection  Memory injection attacks poison the agent’s persistent context, for example its stored plans, past prompts, or retrieved documents, rather than the live user prompt. Patlan et al. (2025a) introduces plan injection where attackers inject malicious steps into the agent’s stored task plan so that, when the Brain component later retrieves that plan at runtime, it unknowingly executes these injected steps alongside the legitimate ones. Similarly, Patlan et al. (2025b) demonstrates how embedding adversarial instructions into an agent’s shared memory store causes the agent to act on those unsafe memories during retrieval-augmented reasoning, triggering unauthorized behaviors without ever altering its immediate input. It is shown by Patlan et al. (2025b) that memory injection achieves greater persistent and stealth compared to prompt injections.
@@
Backdoor Attack A backdoor attack involves injecting a malicious backdoor during the model’s training or fine-tuning phase, so that when a specific trigger phrase or input is later encountered during normal operations, the model executes unintended or harmful behavior (Yang et al., 2024b; Wang et al., 2024; Zhu et al., 2025b).
In particular, we can categorizes backdoor attacks on CUAs into two main forms: (1) The trigger is placed in the agent’s inputs (user queries or environmental observations) (Yang et al., 2024b) and directly manipulates the final output. For example, AgentPoison (Chen et al., 2024b) optimizes textual triggers via mapping poisoned instances to a unique embedding space, ensuring only trigger-containing prompts retrieve the malicious demonstrations. Boisvert et al. fine-tune LLM agents on mixed web and tool-use interaction logs that have been poisoned by inserting a benign-looking <div> with a unique ID into WebArena’s accessibility tree and a #\#EXFILTRATE_\_DATA token into $\tau$-Bench tool sequences, causing agents to execute the hidden malicious actions whenever they encounter these triggers at inference. A RAG backdoor attack embeds malicious payloads and corresponding trigger tokens directly into the agent’s external knowledge documents—such as augmented PDFs or database entries—so that when the agent retrieves those documents during its planning or observation phase, it unwittingly executes the attacker’s instructions without altering its core model parameters. Beyond text, VisualTrap (Ye et al., 2025) poisons GUI grounding data by remaping a tiny, low-salience on-screen mark to specific element–action pairs, causing the agent to execute attacker chosen clicks until the visual trigger appears. Similarly, (Wang et al., 2025f) injects a clean-label backdoor by fine-tuning vision–language mobile agents on a small fraction of screenshots covertly perturbed with an imperceptible visual trigger. More stealthily, Cheng et al. (2025) crafts composite triggers at the goal and interaction levels, using a min–max optimization with supervised contrastive learning and finetuning to ensure benign behavior on clean inputs and precise malicious actions when both trigger conditions are met. (2) The model’s internal reasoning is corrupted without visibly changing the final answer, e.g., covertly calling untrusted APIs (Yang et al., 2024b). Building on the RAG paradigm, Lupinacci et al. (2025) also show that a RAG backdoor attack can simply embed malicious payloads and trigger tokens into external knowledge documents, such as augmented PDFs or database entries, so that during retrieval and planning, the agent unwittingly executes the attacker’s instructions.
Meanwhile, to enhance stealthiness and bypass safety audits, attackers can break the backdoor code into multiple sub-backdoors, each activated by its own distinct trigger phrase or condition. When these sub-backdoors are combined, they enable the model to execute coordinated malicious behaviors (Zhu et al., 2025b). This modular design obscures the overall functionality behind seemingly unrelated trigger fragments, making detection and mitigation significantly more difficult.
6
@@
Chen et al. (2025d) examines how multimodal mobile agents are vulnerable to these attacks. The study shows that when attackers add conflicting or deceptive signals, such as subtle differences in an image combined with misleading text, the agent’s reasoning process struggles to correctly combine the different inputs. As a result, the agent might misinterpret the environment and take the wrong action.
@@
System Sabotage  System sabotage attacks involve manipulating an agent into executing harmful actions that damage the underlying system. In such attacks, adversaries craft inputs to bypass safety mechanisms, causing the agent to perform operations like corrupting memory, damaging critical files, or halting essential processes (Luo et al., 2025b). These attacks are particularly dangerous because they directly target the infrastructure supporting the agent, potentially leading to widespread system failure or irreversible damage.
One example stated in (Luo et al., 2025b) is an attacker requests the agent’s assistance in creating a fork bomb, which is an intentionally crafted command that spawns processes indefinitely and tends to overwhelm the operating system. The user prompt disguises this request as a system “stress test,” persuading the agent to generate code that saturates system resources. Once executed, this fork bomb can cause the OS to become unresponsive or crash.
8